OpenStudy (anonymous):

how do i rewrite sqrt(384)

OpenStudy (anonymous):

i know is sqrt(2^7) x 3 but i forgot how to rewrite that :/

OpenStudy (mathstudent55):

\(\sqrt{384} = \sqrt{2^7 \times 3} = \sqrt{2^6 \times 2 \times 3} = \sqrt{2^6} \times \sqrt{2 \times 3} = 2^3 \sqrt{6} = 8\sqrt{6}\)